OpenStudy (anonymous):

sin 33 degrees = 7/x whats the value of x ? WILL FAN & MEDAL thank you

OpenStudy (anonymous):

ok i'm here

OpenStudy (anonymous):

Yay ! Okay, how do i do this ? @ttb123456789

OpenStudy (anonymous):

12.85 but let me check

OpenStudy (anonymous):

well x=7/(sin33)=7/0.544639035... x=12.85254921

OpenStudy (anonymous):

thank you! @ttb123456789

OpenStudy (anonymous):

no problem tell me the score of your test when you will finish

OpenStudy (anonymous):

Okay, it might be a while though. I'm on question 8 of 25

OpenStudy (anonymous):

mmmm!!

OpenStudy (anonymous):

well tag me i am going to help others

OpenStudy (anonymous):

I made a 92 :D @ttb123456789

OpenStudy (anonymous):

I was wrong?

OpenStudy (anonymous):

@taylor1216

OpenStudy (anonymous):

i have calculated that you had done 2 mistakes.

OpenStudy (anonymous):

I forget which two were wrong